OpenStudy (anonymous):

Given that f(x) = x2 + 3x + 6 and g(x) = , solve for f(g(x)) when x = 1. a. -1 b. 0 c. 3 d. 4 please walk through/explain it, I want to understand it!!

OpenStudy (anonymous):

missing something?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

look carefully at your question i think there is a copy and paste fail

OpenStudy (anonymous):

oh oops its G(x) = (x-3 / 2)

OpenStudy (anonymous):

\[g(x)=\frac{x-3}{2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k step 1 find \[g(1)\]

OpenStudy (anonymous):

you good with that?

OpenStudy (anonymous):

Uh no I think I've learned it a different way than you, I thought you plugged in g(x) into the f(x) equation

OpenStudy (anonymous):

you are being asked for \[f(g(1))\]right?

OpenStudy (anonymous):

yes sorry my bad

OpenStudy (anonymous):

k then you need \(g(1)\) first

OpenStudy (anonymous):

so would it be F(x) = (\frac{ 1-3 }{ 2 })^{2} +3(\frac{ 1-3 }{ 2 } )+ 6\]

OpenStudy (anonymous):

\[ F(x) = (\frac{ 1-3 }{ 2 })^{2} +3(\frac{ 1-3 }{ 2 } )+ 6\] is way way way too much work and although it is correct, not to be rued but it is rather silly lets take it step by step

OpenStudy (anonymous):

\[g(x)=\frac{x-3}{2}, g(1)=\frac{1-3}{2}=-1\] \[f(x)=x^2+3x+6, f(-1)=(-1)^2+3\times (-1)+6=4\]

OpenStudy (anonymous):

So then the answer would be 3???

OpenStudy (anonymous):

i get 4

OpenStudy (anonymous):

\[(-1)^2+3\times (-1)+6=1-3+6=-2+6=4\]

OpenStudy (anonymous):

Oh nevermind I get it now!! Thank you!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now if you had wanted \[f(g(x))\] you would have had to do what you did and write \[f(g(x))=(\frac{ x-3 }{ 2 })^{2} +3(\frac{ x-3 }{ 2 } )+ 6\]

OpenStudy (anonymous):

so technically I was correct

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

@satellite73 okay thank you! do you think you could help me with this other one I'm stuck on?